OpenStudy (anonymous):

The gravitational force exerted on a baseball is 2.19 N down. A pitcher throws the ball horizontally with velocity 16.0 m/s by uniformly accelerating it along a straight horizontal line for a time interval of 151 ms. The ball starts from rest. (a) Through what distance does it move before its release?

OpenStudy (anonymous):

Vf = (Vo=0) +at --> 16 = a*.151 d= (Vo=0)*t + 1/2*(at^2) --> d = 1/2*(16/.151)*(.151)^2